# BioBERT-based Deep Learning and Merged ChemProt-DrugProt for Enhanced Biomedical Relation Extraction

### v1

@@ Abstract
This paper presents a methodology for enhancing relation extraction from biomedical texts, focusing specifically on chemical-gene interactions. Leveraging the BioBERT model and a multi-layer fully connected network architecture, our approach integrates the ChemProt and DrugProt datasets using a novel merging strategy. Through extensive experimentation, we demonstrate significant performance improvements, particularly in CPR groups shared between the datasets. The findings underscore the importance of dataset merging in augmenting sample counts and improving model accuracy. Moreover, the study highlights the potential of automated information extraction in biomedical research and clinical practice.
Introduction
@@
Biomedical literature serves as a vital conduit for various stakeholders within the scientific community, including biomedical researchers, clinicians, and database curators. Through articles, patents, and reports, these individuals disseminate their findings, contributing to the collective knowledge base of the field. However, the sheer volume of literature generated on a daily basis presents a significant challenge, hindering users’ ability to efficiently retrieve relevant information. Consequently, there is a pressing demand for innovative solutions to streamline information retrieval processes. Recognizing this need, Natural Language Processing (NLP) systems have emerged as promising tools to automate the extraction of pertinent information from biomedical texts. By leveraging computational algorithms, NLP systems aim to expedite the identification and extraction of key insights, thereby alleviating the manual burden placed on users. NLP, as a field of study, is dedicated to developing techniques that enable computers to understand and analyze human language in its unstructured form. Within the realm of NLP, Relation Extraction (RE) stands out as a crucial area of focus. RE involves the identification and characterization of relationships between entities mentioned within textual data. By discerning connections between various entities—such as genes, proteins, diseases, and treatments—RE facilitates the extraction of meaningful insights from biomedical literature. This capability holds immense potential for advancing biomedical research, clinical practice, and data curation efforts, ultimately driving innovation and improving outcomes within the healthcare domain.
@@
In the landscape of biomedical literature analysis, a significant portion of existing systems is dedicated to the automatic recognition of mentions pertaining to genes, proteins, and chemicals within textual data. While these systems play a pivotal role in facilitating the identification of individual entities, a noticeable gap exists in their ability to extract and elucidate the intricate interactions between these entities. Indeed, a limited number of approaches have been developed to specifically target the extraction of interactions between genes/proteins and chemicals within textual data. Given this context, there is a clear imperative to delve deeper into the diverse relationships that exist between drugs, chemical compounds, and various biomedical entities, particularly genes and proteins. Systematic extraction of these relationships is essential for enabling comprehensive analysis and exploration of key biomedical properties across a spectrum of applications. By honing in on these interactions, researchers can unravel crucial insights into drug mechanisms, disease pathways, and therapeutic targets, paving the way for advancements in drug discovery, personalized medicine, and disease management. Efforts to enhance the extraction of such relationships hold immense promise for bolstering the efficacy and efficiency of biomedical research and clinical practice. By leveraging advanced Natural Language Processing (NLP) techniques and innovative methodologies, researchers can unlock the full potential of biomedical literature, accelerating the pace of scientific discovery and revolutionizing healthcare delivery.
This paper presents a methodology for improving relation extraction from biomedical texts, focusing on chemical-gene interactions. Leveraging deep learning approaches, specifically the BioBERT model and a fully connected network architecture, the study introduces a merging strategy to integrate the ChemProt and DrugProt datasets. Experimental results demonstrate significant performance enhancements, particularly in CPR groups shared between the datasets. The approach showcases the effectiveness of dataset merging in augmenting sample counts and improving model accuracy. The findings underscore the potential of automated information extraction in biomedical literature analysis and highlight avenues for future research in Natural Language Processing for biomedicine.
@@ 1 Related work
Many approaches have been explored for RE in the biomedical domain and this can be divided into 4 major classes: 1) rule-based, 2) machine learning-based 3) deep learning-based, and 4) contextualized language model-based approaches. Rule-based approaches utilize rules and patterns to identify relations between words. He, et al. [1] proposed a rule-based method to identify relations between chemical reactions synthesis by compiling two dictionaries: the first to identify the entities, and the second to identify the relations between entities that occur in the same sentence. Li, et al. [2] proposed a rule-based method where regular expressions were used to match drug names to a prescription list, followed by the use of co-location information to link the attributes to the drugs. Lowe, et al. [3] used the ChemicalTagger [4], an open-source tool that uses syntactic information to identify relations between entities.
@@
Machine learning-based approaches use machine learning (ML) algorithms and statistical analysis to classify relations. Support vector machines (SVMs) dominated most of the early machine-learning-based approaches [5, 6, 7, 8]. Miller, et al. [9] used token context features, character type features, and semantic features with SVMs. Anick, et al. [10] used lists of n-grams with specific semantics as the features for SVM. Demner-Fushman, et al. [7] incorporated semantic information using concepts from the Unified Medical Language System [11] a repository of biomedical and clinical concepts. They exercised feature reduction through cross-validation. Zhu, et al. [12] focused on revising the learning algorithm by reformulating the SVM into a composite-kernel framework to achieve better performance.
@@
Hybrid approaches combined linguistic pattern matching with ML techniques. Grouin, et al. [6] and Minard, et al. [13] trained an SVM and constructed linguistic patterns manually. They reported that there are advantages of the hybrid approach as linguistic patterns may confirm automatically-induced relations which helps adding confidence to the obtained results. While Yang, et al. [14] applied heuristic rules to generate candidate pairs of possible related entities that were fed into three ML models (SVM, random forest, and gradient boosted trees [15]) to classify the relations. The possible related entities were identified according to their distance from each other as defined by the number of sentence boundaries between the two entities, and multiple classifiers were developed to classify relations based on this distance.
@@
Deep learning is a subfield of machine learning. Sahu, et al. [16] and Luo, et al. [17] utilized convolutional neural network (CNN) to automatically learn feature representations to reduce the need for engineered features. Sahu, et al. [16] applied CNN to build a Sentence-CNN which learns a single sentence-level representation for each relation, using discrete features. Lv, et al. [18] proposed the adoption of a CRF model and applied a deep learning model for feature optimization by the employment of autoencoder and sparsity limitation. Tang, et al. [19] used a hierarchical attention-based convolutional LSTM (ConvLSTM) model to construct a sentence as a multi-dimensional hierarchical sequence, to learn the local and global context information. Wei, et al. [20] proposed a model combining CNN and recurrent neural network (RNN) where they utilized Local context and semantic features as features [21].
@@
Leveraging contextualized language models for RE in the biomedical domain has been gaining attention in the recent times. Alimova, et al. [22] conducted a comparison between three BERT-based models: BERT-uncased [23], BioBERT [24], and Clinical BERT [25] for extracting relations from clinical texts. Copara, et al [26] used a BERT-based method assessing five variations of the BERT language models, including a domain-specific model called ChemBERTa. Mahendran, et al. [27] utilized two general BERT and a BioBERT models to automatically detect relations between chemical compounds/drugs and genes/proteins. Zhang, et al. [28] proposed a hybrid method combining deep learning models with pattern-based rules and built a binary classifier by fine-tuning BioBERT. Mahendran, et al. [29] explored rule-based, deep learning based and BERT-based methods to identify Adverse drug events (ADEs) from clinical text. Mahendran, et al. [30] also proposed combining BERT with graph convolutional network (GCN) to extract information regarding chemicals and chemical reactions from chemical patents.
@@
Zhou, et al. [31] utilized large language models (LLMs) by introducing LEAP, a framework using adaptive instructions and examples to find relationships in clinical data. Yoon, et al. [32] proposed a system that uses LLMs, a combination of weakly labeled data and knowledge bases to achieve better performance than standard methods. Yuan, et al. [33] created a new model, KeBioLM, that uses a well-established knowledge base called UMLS to understand biomedical text by incorporating existing medical knowledge. Tinn, et al. [34] and Delmas, et al. [35] explored training with limited data. Tinn, et al. used a medical vocabulary and specialized pre-training to create robust models for biomedical applications whereas Delmas, et al. explored different techniques including training with synthetic data generated by LLMs. Dunn, et al. [36] presented a simple sequence-to-sequence approach to joint named entity recognition (NER) and RE for complex hierarchical information by leveraging a fine-tuned LLM.
2 Method
2.1 Data
In this study, we employed two datasets, ChemProt and DrugProt, which contain annotations for chemical-gene interactions. These datasets were released as part of the BioCreative challenges  [^1] https://biocreative.bioinformatics.udel.edu/. Both datasets offer manual annotations of chemical and gene entities, along with their relationships, within the context of PubMed abstracts. Notably, DrugProt, released at a later date, exhibits a significantly larger volume of entity mentions and relations compared to ChemProt. This disparity is detailed in Table 1, which presents statistics on the number of abstracts, entities, and relations in each dataset. Both ChemProt and DrugProt categorize their chemical-gene relations into 22 distinct classes. These relations can be further aggregated into ten ChemProt Relation (CPR) groups. The definitions of these ten groups are provided in Table 2, originally sourced from ChemProt’s documentation. To streamline our analysis, we preprocessed both datasets to map all 22 relation categories to these 10 groups, as illustrated in Table LABEL:tab:mapping.
@@ 2.1 Data
Table 2: CPR Group-to-Relations Mapping
CPR Group Relations
CPR:1 PART_OF
CPR:2 REGULATOR, DIRECT_REGULATOR, INDIRECT_REGULATOR
CPR:3 UPREGULATOR, ACTIVATOR, INDIRECT_UPREGULATOR
CPR:4 DOWNREGULATOR, INHIBITOR, INDIRECT_DOWNREGULATOR
CPR:5 AGONIST, AGONIST_ACTIVATOR, AGONIST_INHIBITOR
CPR:6 ANTAGONIST
CPR:7 MODULATOR, MODULATOR_ACTIVATOR, MODULATOR_INHIBITOR
CPR:8 COFACTOR
CPR:9 SUBSTRATE, PRODUCT_OF, SUBSTRATE_PRODUCT_OF
CPR:10 NOT
2.2 Dataset merging methodology
Although the DrugProt and ChemProt datasets share similarities, DrugProt contains a larger number of relations compared to ChemProt. However, due to the presence of overlapping instances in both datasets, straightforward merging is not feasible. To address this challenge, we devised a merging strategy. Initially, we organized entities and relations based on the abstracts in which they appear, resulting in two distinct sets: one comprising all ChemProt abstracts and the other containing all DrugProt abstracts. During the merging process, if an abstract is exclusively present in one set, it is directly added to the merged set. However, in cases where an abstract exists in two versions across both sets, we attempt to merge the entities and relations from both versions. To ensure coherence, any disagreements are disregarded to prevent the introduction of noise into the merged set. With this merging strategy, we found no inconsistencies in terms of text and entity content. However, some relation conflicts were found: 63 relation conflicts were found in the training sets and 7 in the validation sets. For clarity, Table 3. provides the size of the merged dataset and the individual sizes of the ChemProt Relation (CPR) groups within it.
Table 3: Merged ChemProt-DrugProt Dataset
@@ 2.2 Dataset merging methodology
2.3 Relation extraction system
Our Relation Extraction (RE) model comprises two primary components: a BioBERT model and a fully connected top layer referred to as the ”top model.” BioBERT, developed by Lee et al. [24], is a contextual embedding model built upon the architecture of BERT. This model has been pre-trained using a vast corpus of biomedical texts, including PubMed abstracts and full-text articles from PMC. Notably, BioBERT offers two configurations: ’base’ and ’large.’ BioBERT Large’s pre-training was performed on a more extensive vocabulary that encompasses a broader spectrum of biomedical terms compared to BioBERT Base. The key distinctions between these configurations are outlined in the table below:
Table 4: Model Descriptions
BioBERT Base BioBERT Large
Number of Layers 12 24
Number of attention heads per layer 12 16
Number of units in hidden layer 768 1024
Vocabulary Original BERT Original BERT
+ custom 30K biomedical
The top model serves as the classification component in our RE model, and is based on a multi-layer fully connected network. The model begins by processing the embedding of the [CLS] token extracted from BioBERT. This embedded representation then passes through two hidden layers, each consisting of 1024 units, before reaching the output layer. To ensure compatibility with the training data, our output structure’s dimensions are aligned with the count of CPR groups present in the training data. Each entry within the vector of this structure represents the probability of the input being associated with the corresponding CPR group. We have developed two variants of the RE model, each built upon different configurations of BioBERT, as detailed in Table 4. Additionally, to comprehensively assess the system’s performance, we evaluated it both with and without the utilization of the aforementioned top model.
2.4 Experimental details
Each individual model underwent independent training on the designated training dataset, followed by evaluation using the validation dataset. To update the model parameters, we utilized the Adam optimizer with a weight decay of 0.01. We used a variable global learning rate scaling approach in which the learning rate increases linearly during the warmup stage, and is equal to the inverse squared-root of the step count after the warmup stage:

### v2

# Merged ChemProt-DrugProt for Relation Extraction from Biomedical Literature

@@ Abstract
The extraction of chemical-gene relations plays a pivotal role in understanding the intricate interactions between chemical compounds and genes, with significant implications for drug discovery, disease understanding, and biomedical research. This paper presents a data set created by merging the ChemProt and DrugProt datasets to augment sample counts and improve model accuracy. We evaluate the merged dataset using two state of the art relationship extraction algorithms: Bidirectional Encoder Representations from Transformers (BERT) specifically BioBERT, and Graph Convolutional Networks (GCNs) combined with BioBERT. While BioBERT excels at capturing local contexts, it may benefit from incorporating global information essential for understanding chemical-gene interactions. This can be achieved by integrating GCNs with BioBERT to harness both global and local context. Our results show that by integrating the ChemProt and DrugProt datasets, we demonstrated significant improvements in model performance, particularly in CPR groups shared between the datasets. Incorporating the global context using GCN can help increase the overall precision and recall in some of the CPR groups over using just BioBERT.
keywords:
Natural Language Processing , Biomedical Relation Extraction , Graph Convolutional Neural Networks , BioBERT , ChemProt , DrugProt
journal: Journal of Templates
1 Introduction
@@
Biomedical literature serves as a vital conduit for various stakeholders within the scientific community, including biomedical researchers, clinicians, and database curators. Through articles, patents, and reports, these individuals disseminate their findings, contributing to the collective knowledge base of the field. However, the sheer volume of literature generated on a daily basis presents a significant challenge, hindering users’ ability to efficiently retrieve relevant information. Consequently, there is a pressing demand for innovative solutions to streamline information retrieval processes. Recognizing this need, Natural Language Processing (NLP) systems have emerged as promising tools to automate the extraction of pertinent information from biomedical texts. By leveraging computational algorithms, NLP systems aim to expedite the identification and extraction of key insights, thereby alleviating the manual burden placed on users. NLP, as a field of study, is dedicated to developing techniques that enable computers to understand and analyze human language in its unstructured form. Within the realm of NLP, Relation Extraction (RE) stands out as a crucial area of focus. RE involves the identification and characterization of relationships between entities mentioned within textual data. By discerning connections between various entities—such as chemicals, genes, proteins, diseases, and treatments—RE facilitates the extraction of meaningful insights from biomedical literature. This capability holds immense potential for advancing biomedical research, clinical practice, and data curation efforts, ultimately driving innovation and improving outcomes within the healthcare domain.
@@
In the landscape of biomedical literature analysis, a significant portion of existing systems is dedicated to the automatic recognition of mentions pertaining to genes, proteins, and chemicals within textual data. While these systems play a pivotal role in facilitating the identification of individual entities, a noticeable gap exists in their ability to extract and elucidate the intricate interactions between these entities. Indeed, a limited number of approaches have been developed to specifically target the extraction of interactions between genes/proteins and chemicals within textual data. Given this context, there is a clear imperative to delve deeper into the diverse relationships that exist between drugs, chemical compounds, and various biomedical entities, particularly genes and proteins. Systematic extraction of these relationships is essential for enabling comprehensive analysis and exploration of key biomedical properties across a spectrum of applications. By honing in on these interactions, researchers can unravel crucial insights into drug mechanisms, disease pathways, and therapeutic targets, paving the way for advancements in drug discovery, personalized medicine, and disease management. Efforts to enhance the extraction of such relationships hold immense promise for bolstering the efficacy and efficiency of biomedical research and clinical practice. By leveraging advanced NLP techniques and innovative methodologies, researchers can unlock the full potential of biomedical literature, accelerating the pace of scientific discovery and revolutionizing healthcare delivery.
In this work, we focus on the creation of a unified dataset for chemical-gene relation extraction by merging two widely used resources: ChemProt and DrugProt. Although chemical-gene relation extraction has been extensively studied using a variety of deep learning methods—including RNNs [32], LSTMs [27], and CNNs [18]—our aim is to support future research by providing a merged, standardized benchmark dataset that incorporates both local (sentence-level) and global (document-level) contextual information. We evaluate the utility of this dataset by benchmarking two established approaches: a BERT-based model [9] for local context modeling, and GCN-BERT [24], which augments BERT with graph convolutional networks to integrate global context across documents.
Section 3 describes the dataset merging strategy. We assess the quality and effectiveness of the merged dataset using the BERT model in Section 4.1, and evaluate it further with GCN-BERT in Section 4.2. Experimental setup and results are presented in Sections 5 and 6, providing a comparative view of model performance on the new dataset. Section 7 discusses the implications of our findings, and we outline directions for future dataset development and evaluation strategies in Section 8.
@@ 2 Related work
Many approaches have been explored for RE in the biomedical domain. These approaches can be divided into 4 major classes: 1) rule-based, 2) machine learning-based 3) deep learning-based, and 4) contextualized language model-based approaches.
@@
Rule-based approaches utilize rules and patterns to identify relations between words. He, et al. [14] proposed a rule-based method to identify relations between chemical reactions synthesis by compiling two dictionaries, first dictionary they used to identify the entities, and the second to identify the relations between entities that occur in the same sentence. Li, et al. [17] proposed a rule-based method where they used regular expressions to match drug names to a prescription list then co-location information to link the attributes to the drugs. Lowe, et al. [19] used the ChemicalTagger [13] an open-source tool that uses syntactic information to identify relations between entities.
@@
Machine learning-based approaches use machine learning algorithms and statistical analysis to identify traditional features best able to classify relations. SVMs dominated most of the machine-learning-based approaches [29, 12, 8, 33]. Miller, et al. [25] used token context features, character type features, and semantic features with SVMs. Anick, et al. [3] used lists of n-grams with specific semantics as the features for SVM. Demner-Fushman, et al. [8] incorporated semantic information using concepts from the UMLS [4] and exercised feature reduction through cross-validation. Zhu, et al. [46] focused on revising the learning algorithm by reformulating the SVM into a composite-kernel framework to achieve better performance.
@@
Hybrid approaches combined linguistic pattern matching with ML techniques. Grouin, et al. [12] and Minard, et al. [26] also trained an SVM and constructed linguistic patterns manually. They reported that there are advantages of the hybrid approach as linguistic patterns may confirm automatically-induced relations which helps adding confidence to the obtained results. While Yang, et al. [39] applied heuristic rules to generate candidate pairs of possible related entities that were fed into three ML models (SVM, RF, and Gradient Boosting [11]) to classify the relations. The possible related entities were identified according to their distance from each other as defined by the number of sentence boundaries between the two entities, and multiple classifiers were developed to classify relations based on this distance.
@@
Deep learning is a field derived from machine learning. Sahu, et al. [31] and Luo, et al. [20] utilized convolutional neural network (CNN) to automatically learn feature representations to reduce the need for engineered features. Sahu, et al. [31] applied CNN to build a Sentence-CNN which learns a single sentence-level representation for each relation, using discrete features. Lv, et al. [21] proposed the adoption of a CRF model and applied a deep learning model for features optimization by the employment of autoencoder and sparsity limitation. Tang, et al. [36] used a hierarchical attention-based convolutional LSTM (ConvLSTM) model to construct a sentence as a multi-dimensional hierarchical sequence, to learn the local and global context information. Wei, et al. [38] proposed a model combining CNN and RNN where they utilized local context and semantic features as features [30].
@@
Leveraging contextualized language models for RE in the biomedical domain has been gaining attention. Alimova, et al. [1] conducted a comparison between three BERT-based models: BERT-uncased [9], BioBERT [16], and Clinical BERT [2] for extracting relations from clinical texts. Copara, et al [6] used a BERT-based method assessing five variations of the BERT language models, including a domain-specific model called ChemBERTa. Zhang, et al. [43] proposed a hybrid method combining deep learning models with pattern-based rules and built a binary classifier by fine-tuning BioBERT. In previous works, we utilized two general BERT and a BioBERT models to automatically detect relations between chemical compounds/drugs and genes/proteins [23], and explored rule-based, deep learning based and BERT-based methods to identify adverse drug events (ADEs) from clinical text [22].
Several groups have explored combining BERT-based models with graph convolutional networks (GCNs) for biomedical NLP tasks. Lai & Lu [15] presented an approach that combine BERT and a graph transformer with neighbor attention to capture relations across sentence boundaries. Our previous work proposed combining BERT with Graph Convolutional Neural Networks to extract relationships between chemicals reactions and their parameters from chemical patents [24]. Subsequently, a number of other systems have been proposed. For example, Nojoo Kambar et al. [28] proposed a model that integrates a BERT-based encoder with a GCN for joint entity and relation extraction of Genes and Chemicals from biomedical texts. While their model is specific to Gene and Chemicals, our previous model is more generalizable across a variety of different entity types.
@@
Zhou, et al. [45] utilize large language models (LLMs) by introducing LEAP, a framework using adaptive instructions and examples to find relationships in clinical data. Yoon, et al. [40] proposed a system that uses LLMs, a combination of weakly labeled data and knowledge bases to achieve better performance than standard methods. Yuan, et al. [41] created a new model, KeBioLM, that uses a well-established knowledge base called UMLS to understand biomedical text by incorporating existing medical knowledge. Tinn, et al. [37] and Delmas, et al. [7] explored training with limited data. Tinn, et al. used a medical vocabulary and specialized pre-training to create robust models for biomedical applications whereas Delmas, et al. explored different techniques including training with synthetic data generated by LLMs. Dunn, et al. [10] presented a simple sequence-to-sequence approach to joint NER and RE for complex hierarchical information by leveraging a fine-tuned LLM.
3 Data
In this section, we describe the ChemProt and DrugProt datasets, our merging strategy and resulting new, larger and more comprehensive dataset.
3.1 Datasets
The ChemProt and DrugProt datasets, both released as part of the BioCreative challenge  [^1] https://biocreative.bioinformatics.udel.edu/, contain chemical-protein/gene interactions and provide manual annotations for these entities along with their relations within their respective PubMed abstracts. There are differences between the two datasets in terms of the number of abstracts, entities, and relations. Notably, due to its later release, DrugProt contains more chemical-gene entity relations compared to ChemProt, as indicated in Table 1.
@@ 3.1 Datasets
Figure 1 illustrates sample sentences with annotated entities and their relations.
Figure 1: Illustration of entities and relations in from sample input sentences
Both ChemProt and DrugProt classify their chemical-gene relations into 22 categories, further organized into 10 ChemProt Relation (CPR) groups. The definitions of these groups, as outlined in ChemProt’s documentation, guided our preprocessing efforts. We systematically mapped all 22 relation categories to these 10 groups, as shown below, to facilitate a more streamlined analysis:
  • 1.
    CPR:1 - PART_OF
  • 2.
    CPR:2 - REGULATOR, DIRECT_REGULATOR, INDIRECT_REGULATOR
  • 3.
    CPR:3 - UPREGULATOR, ACTIVATOR, INDIRECT_UPREGULATOR
  • 4.
    CPR:4 - DOWNREGULATOR, INHIBITOR, INDIRECT_DOWNREGULATOR
  • 5.
    CPR:5 - AGONIST, AGONIST_ACTIVATOR, AGONIST_INHIBITOR
  • 6.
    CPR:6 - ANTAGONIST
  • 7.
    CPR:7 - MODULATOR, MODULATOR_ACTIVATOR, MODULATOR_INHIBITOR
  • 8.
    CPR:8 - COFACTOR
  • 9.
    CPR:9 - SUBSTRATE, PRODUCT_OF, SUBSTRATE_PRODUCT_OF
  • 10.
    CPR:10 - NO RELATION
3.2 Dataset Merging Processing
We developed a merging strategy to combine the ChemProt and DrugProt datasets due to the considerable overlap between the two. As DrugProt contains a larger number of relations than ChemProt, straightforward merging was challenging.
We initially organized entities and relations based on the abstracts in which they appear, resulting in distinct sets for ChemProt and DrugProt abstracts. During the merging process, abstracts exclusively present in one set are directly added to the merged set. In instances where an abstract exists in both sets, we merged entities and relations from both datasets for that abstract. For these common entities, no inconsistencies were detected in terms of text and entity content; however, relation conflicts were identified. In the training sets, 63 relation conflicts were identified, along with 7 conflicts in the validation sets. These conflicts were resolved manually. The size of the merged dataset and the individual ChemProt Relation (CPR) group sizes are detailed in Table 2.
@@ 3.2 Dataset Merging Processing
Table 2: Merged ChemProt-DrugProt Dataset
The BRAT rapid annotation tool [34] has been widely used in the NLP research community for various annotation tasks, including named entity recognition, relation extraction, and co-reference resolution. This annotation format is designed in particular for structured annotation, where the notes are not freeform text but have a fixed form that can be automatically processed and interpreted by a computer. All annotations follow the same basic structure: Each line contains one annotation, and each annotation is given an ID that appears first on the line, separated from the rest of the annotation by a single TAB character. The rest of the structure varies by annotation type. We converted the merged dataset from JSON to annotations in BRAT format, utilizing the .ann (annotated files) and .txt files for each abstract in the dataset.
3.3 Training Data
The ChemProt and DrugProt datasets were part of the BioCreative challenge, thus only training and validation data were made available, i.e, the test data was not made available. We created merged ChemProt-DrugProt training and validation datasets as described above. For model training, the merged validation dataset was used as the held-out test set for final evaluation. The merged train dataset was partitioned using an 80/20 ratio with stratified sampling to create train and validation datasets for model development and hyperparameter tuning.
For our reported results, CPR groups 7 and 8 were omitted due to their insufficient representation. Since the sample counts in these groups are so small compared to the other groups, the model may never learn these classes, leading to poor generalization. Additionally, to augment the negative class (CPR:10), we introduced additional instances by selecting pairs of chemical and gene entities lacking any association in each input sentence. In other words, if a given chemical-gene pair did not have any annotated relation, they were assigned a CPR-10 (No Relation) label. This approach generated more explicit negative samples to prevent the model from making spurious relation predictions where there was none.
4 Relation Extraction Methods
4.1 BioBERT Model Setup
Our BioBERT-based model comprises of two primary components: a BioBERT model and a fully connected top layer referred to as the ”top model.” BioBERT, developed by Lee et al. [16], is a contextual embedding model built upon the architecture of BERT. This model has been pre-trained using a vast corpus of biomedical texts, including PubMed abstracts and full-text articles from PMC. The top model is a multi-layer fully connected network that is stacked on top of BioBERT and serves as the classification component in our RE model. The model begins by processing the embedding of the [CLS] token extracted from BioBERT. This embedded representation then passes through two hidden layers, each consisting of 1024 units, before reaching the output layer. To ensure compatibility with the training data, our output structure’s dimensions are aligned with the count of CPR groups present in the training data. Each entry within the vector of this structure represents the probability of the input being associated with the corresponding CPR group.
The BioBERT model was fine-tuned on the training dataset, followed by evaluation using the validation dataset. To update the model parameters, we utilized the Adam optimizer with a weight decay of 0.01. We used a variable global learning rate scaling approach in which the learning rate increases linearly during the warmup stage, and is equal to the inverse squared-root of the step count after the warmup stage: